OpenStudy (anonymous):

2x^(2)-7x-4/x^(2) * 8x^(2)-28x+12/4x^(2)-1 Multiply

OpenStudy (mertsj):

Factor and divide out the common factors before completing the multiplication.

OpenStudy (mertsj):

\[\frac{2x^2-7x-4}{x^2}\times \frac{8x^2-28x+12}{4x^2-1}\]

OpenStudy (mertsj):

First factor out the 4

OpenStudy (mertsj):

4(2x^2-7x+3)

OpenStudy (anonymous):

ok....

OpenStudy (mertsj):

Now factor 2x^2-7x+3

OpenStudy (anonymous):

\[\frac{ (2x+1)(x-4)}{ x^2 }*\frac{ (4(2x-1)(x-3) }{ (2x-1)(2x+1) }\]

OpenStudy (mertsj):

Can you factor or not?

OpenStudy (anonymous):